### v3

# VMamba: Visual State Space Model

Abstract
Designing computationally efficient network architectures persists as an ongoing necessity in computer vision. In this paper, we transplant Mamba, a state-space language model, into VMamba, a vision backbone that works in linear time complexity. At the core of VMamba lies a stack of Visual State-Space (VSS) blocks with the 2D Selective Scan (SS2D) module. By traversing along four scanning routes, SS2D helps bridge the gap between the ordered nature of 1D selective scan and the non-sequential structure of 2D vision data, which facilitates the gathering of contextual information from various sources and perspectives. Based on the VSS blocks, we develop a family of VMamba architectures and accelerate them through a succession of architectural and implementation enhancements. Extensive experiments showcase VMamba’s promising performance across diverse visual perception tasks, highlighting its advantages in input scaling efficiency compared to existing benchmark models. Source code is available at https://github.com/MzeroMiko/VMamba.
1 Introduction
Visual representation learning stands as a fundamental research area in computer vision, which has witnessed remarkable progress in the era of deep learning. To represent complex patterns in vision data, two primary categories of backbone networks, i.e., Convolution Neural Networks (CNNs) [49, 28, 30, 54, 38] and Vision Transformers (ViTs) [13, 37, 58, 68], have been proposed and extensively utilized in a variety of visual tasks. Compared to CNNs, ViTs generally demonstrate superior learning capabilities on large-scale data due to the integration of the self-attention mechanism [59, 13]. However, the quadratic complexity of self-attention w.r.t. the number of tokens introduces substantial computational overhead in downstream tasks involving large spatial resolutions.
To tackle this challenge, considerable efforts have been made to enhance the efficiency of attention computation [55, 37, 12]. However, existing approaches either impose limitations on the size of the effective receptive field [37] or experience evident performance degradation across diverse tasks [31, 62]. This motivates us to develop a novel architecture for vision data, preserving the inherent advantages of the vanilla self-attention mechanism, i.e., global receptive fields and dynamic weighting parameters [23].
Recently, Mamba [17], a novel State Space Model (SSM) [17, 44, 61] in the field natural language processing (NLP), has emerged as a highly promising approach for long sequence modeling with linear complexity. Drawing inspiration from this advancement, we introduce VMamba, a vision backbone integrating SSM-based blocks to facilitate efficient visual representation learning. However, the core algorithm of Mamba, i.e., the parallelized selective scan operation, is essentially designed for processing one-dimensional sequential data. This poses a challenge when attempting to adapt it for processing vision data, which inherently lacks a sequential arrangement of visual components. To address this issue, we propose 2D Selective Scan (SS2D), a four-way scanning mechanism tailored for spatial domain traversal. In contrast to the self-attention mechanism (Figure 1 (a)), SS2D ensures that each image patch gains contextual knowledge exclusively through a compressed hidden state computed along the corresponding scanning path (Figure 1 (b)), thereby reducing the computational complexity from quadratic to linear.
Figure 1: Comparison of correlation establishment between image patches via (a) self-attention and (b) the proposed 2D-Selective-Scan (SS2D). Red boxes indicate the query image patch, with patch opacity representing the degree of information loss.
Upon the VSS blocks, we develop a family of VMamba architectures (i.e., VMamba-Tiny/Small/Base) and accelerate them through a series of architectural enhancements and implementation optimizations. Compared to benchmark vision models built on CNN (ConvNeXt [38]), ViT (Swin [37] and HiViT [68]), and SSM (S4ND [45] and Vim [71]), VMamba consistently achieves superior image classification accuracy on ImageNet-1K [9] across model scales. Specifically, VMamba-Base achieves a top-1 accuracy of $83.9\%$, surpassing Swin by $+0.4\%$, with a throughput exceeding that of Swin by a substantial margin over $40\%$ ($646$ vs. $458$). The superiority of VMamba extends across various downstream tasks, with VMamba-Tiny/Small/Base achieving $47.3\%$/$48.7\%$/$49.2\%$ mAP in object detection on COCO [34] (1×1$\times$1 × training schedule). This outperforms Swin by $4.6\%$/$3.9\%$/$2.3\%$ and ConvNeXt by $3.1\%$/$3.3\%$/$2.2\%$, respectively. As for single-scale semantic segmentation on ADE20K [70], VMamba-Tiny/Small/Base achieves $47.9\%$/$50.6\%$/$51.0\%$ mIoU, which surpasses Swin by $3.4\%$/$3.0\%$/$2.9\%$ and ConvNeXt by $1.9\%$/$1.9\%$/$1.9\%$, respectively. Furthermore, unlike ViT-based models, which experience quadratic growth in computational complexity with the number of input tokens, VMamba exhibits linear growth in FLOPs while maintaining comparable performance. This underscores its state-of-the-art input scalability.
The contributions of this study are summarized as follows:
  • We propose VMamba, an SSM-based vision backbone network for visual representation learning with linear time complexity. A series of improvements in architectural design and implementation details are adopted to improve the inference speed of VMamba.
  • We introduce 2D Selective Scan (SS2D) to bridge the gap between 1D array scanning and 2D plane traversal, facilitating the extension of selective SSM to process vision data.
  • Without bells and whistles, VMamba demonstrates promising performance across a range of visual tasks, including image classification, object detection, and semantic segmentation. It also exhibits remarkable adaptability w.r.t. the length of the input sequence, showcasing linear growth in computational complexity.
2 Related Work
Convolutional Neural Networks (CNNs).
Starting from AlexNet [32], significant efforts have been dedicated to enhancing the modeling capability [49, 53, 28, 30] and computational efficiency [29, 54, 66, 47] of CNN-based models across a spectrum of visual tasks. More sophisticated operators, such as the depth-wise convolution [29] and deformable convolution [5, 72], have been proposed for increasing the flexibility and efficacy of CNNs. Recently, drawing inspiration from the success of Transformers [60], modern CNNs [38] have demonstrated promising performance by incorporating long-range dependencies [11, 48, 35] and dynamic weights [24] into their architectural design.
Vision Transformers (ViTs).
As the most representative pioneering work, ViT [13] explores the effectiveness of vision models based on vanilla Transformer architecture and reveals the significance of large-scale pre-training in improving the performance in image classification. To mitigate ViT’s reliance on extremely large datasets, DeiT [58] introduces a teacher-student distillation strategy to transfer knowledge encapsulated in CNN models to ViTs, showcasing the importance of the inductive bias in visual perception. Following this line of thought, subsequent studies propose hierarchical ViTs [37, 12, 63, 40, 68, 56, 6, 10, 69, 1].
Another line of research focuses specifically on enhancing the computational efficiency of the self-attention mechanism, which serves as the cornerstone of ViTs. By formulating self-attention as the linear dot-product of kernel feature maps, Linear Attention [31] leverages the associativity property of matrix products to reduce the computational complexity from quadratic to linear. GLA [67] further introduces a hardware-efficient variant of linear attention, which balances memory movement against parallelizability. RWKV [46] also leverages the linear attention mechanism to combine the parallelizable training of transformers with the efficient inference of recurrent neural networks (RNNs). RetNet [52] integrates an additional gating mechanism to enable a parallelizable computation path as an alternative to recurrence, and RMT [15] extends it for visual representation learning by incorporating the temporal decay mechanism into the spatial domain.
State Space Models (SSMs).
Despite the widespread adoption of ViT architectures in vision tasks, the quadratic complexity of self-attention presents a significant challenge when dealing with long input sequences, (e.g., high-resolution images). Among the attempts to improve the scaling efficiency [8, 7, 46, 52, 42], SSMs have emerged as compelling alternatives to Transformers, attracting significant attention from the community. Gu et al. [21] demonstrated the potential of SSM-based models in handling the long-range dependency with the HiPPO initialization [18]. To improve the practical feasibility, S4 [20] proposed to normalize the parameter matrices into a diagonal structure. Subsequently, various structured SSM models have emerged, each introducing different architectural enhancements. These include models with complex-diagonal structures [22, 19], support for multiple-input multiple-output [50], decomposition of diagonal plus low-rank operations [25], and selection mechanisms [17]. These advancements have also been integrated into larger representation models [44, 42, 16], further demonstrating the versatility and scalability of structured state space models in various applications. While these models primarily focus on applying SSMs to long-range and sequential data like text and speech, there has been limited exploration of SSMs’ application to vision data with two-dimensional structures.
3 Preliminaries
Formulation of SSMs.
Originating from the Kalman filter [33], SSMs can be regarded as linear time-invariant (LTI) systems that map the input stimulation $u(t)\in\mathbb{R}$ to response $y(t)\in\mathbb{R}$ through the hidden state $\mathbf{h}(t)\in\mathbb{R}^{N}$. Concretely, continuous-time SSMs can be formulated as linear ordinary differential equations (ODEs) as follows,
$\begin{split}\mathbf{h^{\prime}}(t)&=\mathbf{A}\mathbf{h}(t)+\mathbf{B}u(t),\\ y(t)&=\mathbf{C}\mathbf{h}(t)+Du(t),\end{split}$ (1)
where $\mathbf{A}\in\mathbb{R}^{N\times N}$, $\mathbf{B}\in\mathbb{R}^{N\times 1}$, $\mathbf{C}\in\mathbb{R}^{1\times N}$, and $D\in\mathbb{R}^{1}$ are the weighting parameters.
Discretization of SSM.
To be integrated into deep models, continuous-time SSMs must undergo discretization in advance. Concretely, consider the time interval $[t_{a},t_{b}],$the analytic solution to the hidden state variable $\mathbf{h}(t)|_{t={t_{b}}}$ can be expressed as
[…]
By sampling with the time-scale parameter $\boldsymbol{\Delta}$ (i.e., $d\tau|_{t_{i}}^{t_{i+1}}=\Delta_{i}$), $h(t_{b})$ can be discretized by
$\mathbf{h}_{b}=e^{\mathbf{A}(\Delta_{a}+...+\Delta_{b-1})}\left(\mathbf{h}_{a}+\sum^{b-1}_{i=a}\mathbf{B}_{i}u_{i}e^{-\mathbf{A}(\Delta_{a}+...+\Delta_{i})}\Delta_{i}\right),$ (3)
where $[a,b]$ is the corresponding discrete step interval. Notably, this formulation approximates the result obtained by the zero-order hold (ZOH) method, which is frequently utilized in the literature of SSM-based models (please refer to Appendix A for detailed proof).
Selective Scan Mechanism.
To tackle the limitation of LTI SSMs (Eq. 1) in capturing the contextual information, Gu et al. [17] propose a novel parameterization method for SSMs that integrates an input-dependent selection mechanism (referred to as S6). However, in the case of selective SSMs, the time-varying weighting parameters present a challenge for efficient computation of hidden states, as convolutions do not accommodate dynamic weights and are consequently rendered inapplicable. Nevertheless, as the recurrence relation of $h_{b}$ in Eq. 3 can be derived, the response $y_{b}$ can still be efficiently computed using associative scan algorithms [2, 43, 51] with linear complexity (the detailed explanation is deferred to Appendix B).
4 VMamba: Visual State Space Model
4.1 Network Architecture
We develop VMamba at three scales: VMamba-Tiny, VMamba-Small, and VMamba-Base (referred to as VMamba-T, VMamba-S, and VMamba-B, respectively). An overview of the architecture of VMamba-T is illustrated in Figure 3 (a), and detailed configurations are provided in Appendix  E. The input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ is first partitioned into patches by a stem module, resulting in a 2D feature map with the spatial dimension of $H/4\times W/4$. Subsequently, multiple network stages are employed to create hierarchical representations with resolutions of $H/8\times W/8$, $H/16\times W/16$, and $H/32\times W/32$. Each stage comprises a down-sampling layer (except for the first stage) followed by a stack of Visual State Space (VSS) blocks.
The VSS blocks serve as the visual counterpart to Mamba blocks [17] (Figure 3 (b)) for representation learning. The initial architecture of VSS blocks (referred to as the ‘vanilla VSS Block’ in Figure 3 (c)) is formulated by substituting the S6 module, which stands as the core of Mamba in concurrently achieving global receptive fields, dynamic weights (i.e., selectivity), and linear complexity, with the newly proposed 2D-Selective-Scan (SS2D) module (to be introduced in the following subsection). To further enhance computational efficiency, we eliminate the entire multiplicative branch (encircled by the red box in Figure 3 (c)), as the effect of the gating mechanism has been achieved by the selectivity of SS2D. Consequently, the resulting VSS block (depicted in Figure 3 (d)) consists of a single network branch with two residual modules, mimicking the architecture of a vanilla Transformer block [60]. Throughout this paper, all results are obtained using VMamba models built with VSS blocks in this architecture.
Figure 2: Illustration of 2D-Selective-Scan (SS2D). Input patches are traversed along four different scanning paths (Cross-Scan), and each sequence is independently processed by distinct S6 blocks. Subsequently, the results are merged to construct a 2D feature map as the final output (Cross-Merge).
Figure 3: Left: Illustration of (a) the overall architecture of VMamba, and (b) - (d) the structure of Mamba and VSS blocks. Right: Comparison of variants of VMamba and benchmark methods in classification accuracy and computational efficiency.
4.2 2D-Selective-Scan for Vision Data (SS2D)
While the sequential nature of the scanning operation in S6 aligns well with NLP tasks involving temporal data, it poses a significant challenge when applied to vision data, which is inherently non-sequential and encompasses spatial information (e.g., local texture and global structure). To address this issue, S4ND [45] reformulates SSM with convolutional operations, directly extending the kernel from 1D to 2D through outer-product. However, such modification prevents the weights from being input-independent, resulting in a limited capacity for capturing contextual information. Therefore, we adhere to the selective scan approach [17] for input processing, and propose the 2D-Selective-Scan (SS2D) module to adapt S6 to vision data without compromising its advantages.
As illustrated in Figure 2, data forwarding in SS2D involves three steps: cross-scan, selective scanning with S6 blocks, and cross-merge. Given the input data, SS2D first unfolds input patches into sequences along four distinct traversal paths (i.e., Cross-Scan), processes each patch sequence using a separate S6 block in parallel, and subsequently reshapes and merges the resultant sequences to form the output map (i.e., Cross-Merge). By adopting complementary 1D traversal paths, SS2D enables each pixel in the image to effectively integrate information from all other pixels in different directions, thereby facilitating the establishment of global receptive fields in the 2D space.
4.3 Accelerating VMamba
As shown in Figure 3 (e), the VMamba-T model with vanilla VSS blocks (referred to as ‘Vanilla VMamba’) achieves a throughput of $426$ images/s and consists of $22.9$M parameters with $5.6$G FLOPs. Despite the state-of-the-art $82.2\%$ classification accuracy (in the tiny-level, outperforms Swin-T [37] by $0.9\%$), the low throughput and high memory overhead pose significant challenges for the practical deployment of VMamba.
[…]
  1. Step (a)
    ($+0.0\%$, $+41$ img/s) by re-implementing Cross-Scan and Cross-Merge in Triton.
  2. Step (b)
    ($+0.0\%$, $-3$ img/s) by adjusting the CUDA implementation of the selective scan to accommodate float16 input and float32 output. This remarkably enhances the training efficiency (throughput from $165$ to $184$), despite slight speed fluctuation at test time.
  3. Step (c)
    ($+0.0\%$, $+174$ img/s) by substituting the relatively slow einsum in selective scan with a linear transformation (i.e., torch.nn.functional.linear). We also adopt the tensor layout of (B, C, H, W) to eliminate unnecessary data permutations.
  4. Step (d)
    ($-0.6\%$, $+175$ img/s) by introducing MLP into VMamba due to its computational efficiency. We also discard the DWConv (depth-wise convolutional [24]) layers and change the layer configuration from [2,2,9,2] to [2,2,2,2] to lower FLOPs.
  5. Step (e)
    ($+0.6\%$, $+366$ img/s) by discarding the entire multiplicative branch, adopting the VSS block illustrate in Figure 3 (d), and reducing the parameter ssm-ratio (the feature expansion factor) from $2.0$ to $1.0$. This allows us to raise the layer numbers to [2,2,5,2] while reducing the FLOPs.
  6. Step (f)
    ($+0.3\%$, $+161$ img/s) by reducing the parameter d_state (the SSM state dimension) from $16.0$ to $1.0$. This allows us to raise ssm-ratio back to $2.0$ and introduce the DWConv layers without increasing the FLOPs.
  7. Step (g)
    ($+0.1\%$, $+346$ img/s) by reducing the ssm-ratio to $1.0$ while changing the layer configuration from [2,2,5,2] to [2,2,8,2].
5 Experiments
In this section, we conduct a series of experiments to evaluate the performance of VMamba and compare it with popular benchmark models across various visual tasks. We also validate the effectiveness of the proposed 2D feature map traversing method by comparing it to alternative approaches. Additionally, we analyze the characteristics of VMamba by visualizing its effective receptive field (ERF) and activation map, and examining its scalability with increasingly long input sequences. We primarily followed the hyper-parameter settings and experimental configurations of Swin [37]. Please refer to Appendix E and F for detailed experiment settings, and Appendix H for more ablations. All experiments are conducted on a server with 8 $\times$ NVIDIA Tesla-A100 GPUs.
5.1 Image Classification
We evaluate VMamba’s performance in image classification on ImageNet-1K [9], and the comparison results against benchmark methods are summarized in Table 1. With similar FLOPs, VMamba-T achieves a top-1 accuracy of $82.6\%$, outperforming DeiT-S by $2.8\%$ and Swin-T by $1.3\%$. Notably, VMamba maintains its performance advantage at both Small and Base scale. For instance, VMamba-B achieves a top-1 accuracy of $83.9\%$, surpassing DeiT-B by $2.1\%$ and Swin-B by $0.4\%$.
In terms of computational efficiency, VMamba-T achieves a throughput of 1,686 images/s, which is either better or comparable to state-of-the-art methods. This advantage persists across VMamba-S and VMamba-B, with a throughput of $877$ images/s and $646$ images/s, respectively. Compared to SSM-based models, the throughput of VMamba-T is 1.47×1.47$\times$1.47 × higher than S4ND-Conv-T [45] and 1.08×1.08$\times$1.08 × higher than Vim-S [71], while maintaining a performance lead over them by a clear margin of $0.4\%$ and $2.1\%$, respectively.
Table 1: Performance comparison on ImageNet-1K. Throughput values are measured with an A100 GPU and an AMD EPYC 7542 CPU, using the toolkit released by [64], following the protocol proposed in [37]. All images are of size $224\times 224$.
  Model Params FLOPs TP. Top-1
(M) (G) (img/s) (%)
Transformer-Based
DeiT-S [58] 22M 4.6G 1761 79.8
DeiT-B [58] 86M 17.5G 503 81.8
HiViT-T [68] 19M 4.6G 1393 82.1
HiViT-S [68] 38M 9.1G 712 83.5
HiViT-B [68] 66M 15.9G 456 83.8
Swin-T [37] 28M 4.5G 1244 81.3
Swin-S [37] 50M 8.7G 718 83.0
Swin-B [37] 88M 15.4G 458 83.5
XCiT-S24 [1] 48M 9.2G 671 82.6
XCiT-M24 [1] 84M 16.2G 423 82.7
  Model Params FLOPs TP. Top-1
(M) (G) (img/s) (%)
ConvNet-Based
ConvNeXt-T [38] 29M 4.5G 1198 82.1
ConvNeXt-S [38] 50M 8.7G 684 83.1
ConvNeXt-B [38] 89M 15.4G 436 83.8
SSM-Based
S4ND-Conv-T [45] 30M - 683 82.2
S4ND-ViT-B [45] 89M - 397 80.4
Vim-S [71] 26M - 811 80.5
VMamba-T 30M 4.9G 1686 82.6
VMamba-S 50M 8.7G 877 83.6
VMamba-B 89M 15.4G 646 83.9
5.2 Downstream Tasks
In this sub-section, we assess the performance of VMamba on downstream tasks, including object detection and instance segmentation on MSCOCO2017 [34], and semantic segmentation on ADE20K [70]. The training framework is developed based on the MMDetection [3] and MMSegmenation [4] library, and we follow [36] to use the Mask-RCNN [27] and UperNet [65] as the detection and segmentation network respectively.
Table 2: Left: Results of object detection and instance segmentation on MSCOCO. $AP^{b}$ and $AP^{m}$ denote box AP and mask AP, respectively. FLOPs are calculated with input size $1280\times 800$. The notation ‘1×1$\times$1 ×’ refers to models fine-tuned for $12$ epochs, and ‘3×3$\times$3 ×MS’ denotes multi-scale training for $36$ epochs. Right: Results of semantic segmentation on ADE20K. FLOPs are calculated with an input size of $512\times 2048$. ‘SS’ and ‘MS’ denote single-scale and multi-scale testing, respectively.
                 Mask R-CNN 1$\times$ schedule
[…]
Swin-T 42.7 39.3 48M 267G
ConvNeXt-T 44.2 40.1 48M 262G
VMamba-T 47.3 42.7 50M 271G
Swin-S 44.8 40.9 69M 354G
ConvNeXt-S 45.4 41.8 70M 348G
VMamba-S 48.7 43.7 70M 349G
Swin-B 46.9 42.3 107M 496G
ConvNeXt-B 47.0 42.7 108M 486G
VMamba-B 49.2 44.1 108M 485G
Mask R-CNN 3$\times$ MS schedule
Swin-T 46.0 41.6 48M 267G
ConvNeXt-T 46.2 41.7 48M 262G
NAT-T 47.7 42.6 48M 258G
VMamba-T 48.8 43.7 50M 271G
Swin-S 48.2 43.2 69M 354G
ConvNeXt-S 47.9 42.9 70M 348G
NAT-S 48.4 43.2 70M 330G
VMamba-S 49.9 44.2 70M 349G
                 ADE20K with crop size 512
Backbone
mIOU
mIOU
Params FLOPs
ResNet-50 42.1 42.8 67M 953G
DeiT-S + MLN 43.8 45.1 58M 1217G
Swin-T 44.5 45.8 60M 945G
ConvNeXt-T 46.0 46.7 60M 939G
NAT-T 47.1 48.4 58M 934G
Vim-S 44.9 - 46M -
VMamba-T 47.9 48.8 62M 949G
ResNet-101 43.8 44.9 86M 1030G
DeiT-B + MLN 45.5 47.2 144M 2007G
Swin-S 47.6 49.5 81M 1039G
ConvNeXt-S 48.7 49.6 82M 1027G
NAT-S 48.0 49.5 82M 1010G
VMamba-S 50.6 51.2 82M 1028G
Swin-B 48.1 49.7 121M 1188G
ConvNeXt-B 49.1 49.9 122M 1170G
NAT-B 48.5 49.7 123M 1137G
RepLKNet-31B 49.9 50.6 112M 1170G
VMamba-B 51.0 51.6 122M 1170G
Object Detection and Instance Segmentation.
[…]
Semantic Segmentation.
Consistent with previous experiments, VMamba demonstrates superior performance in semantic segmentation on ADE20K with a comparable amount of parameters. As shown in Table 2, VMamba-T is $3.4\%$ mIoU higher than Swin-T and $1.9\%$ higher than ConvNeXt-T in the Single-Scale (SS) setting, and the advantage persists with Multi-Scale (MS) input. For models at the Small and Base level, VMamba-S/B outperforms NAT-S/B [26] by $2.6\%$/$2.5\%$ mIoU under the SS setting, and $1.7\%$/$1.9\%$ mIoU under the MS setting.
Figure 4: Illustration of the adaptability of VMamba to (a) downstream tasks, and (b) input image with progressively increased resolutions. Swin-T denotes Swin-T tested with scaled window sizes.
Discussion
Experiment results in this subsection demonstrate VMamba’s adaptability to object detection, instance segmentation, and semantic segmentation. In Figure 4 (a), we compare VMamba’s performance with Swin and ConvNeXt, highlighting its advantages in addressing downstream tasks with comparable classification accuracy on ImageNet-1K. VMamba also exhibits greater tolerance to changes in input resolution, with linear growth in FLOPs (see Figure 4 (b)) and modest accuracy drop (further discussed in Appendix G), making it more effective and efficient compared to ViT-based methods when adapting to downstream tasks with inputs of larger spatial resolutions. This aligns with Mamba’s advanced capability in efficient long sequence modeling [17].
5.3 Analysis and Discussion
Figure 5: Illustration of the activation map for query patches indicated by red stars. Visualization results in (b) and (c) are obtained by combining the activation map of each scanning path in SS2D.
Relationship between SS2D and Self-Attention.
[…]
$\mathbf{Y}^{(j)}=\left[\mathbf{Q}\odot\mathbf{w}^{(j)}\right]\mathbf{h_{a}}^{(j)}+\left[\left(\mathbf{Q}\odot\mathbf{w}^{(j)}\right)\left(\frac{\mathbf{K}}{\mathbf{w}^{(j)}}\right)^{\top}\odot\mathbf{M}\right]\mathbf{V}^{(j)},$ (4)
where $\mathbf{h_{a}}\in\mathbb{R}^{D_{k}}$ is the hidden state at step $a$, $\mathbf{M}$ denotes the temporal mask matrix of size $T\times T$ with the lower triangular part set to 1 and elsewhere 0, and $\odot$ denotes element-wise product. Detailed derivations are deferred to Appendix C.
In Eq. 4, the matrix multiplication process involving $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ closely resembles the mechanism of self-attention, despite the inclusion of $\mathbf{w}$. Concretely, the formulation of each element in $\mathbf{w}\coloneqq\left[\mathbf{w}_{1};\ldots;\mathbf{w}_{T}\right]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, i.e., $\mathbf{w}_{i}\in\mathbb{R}^{D_{k}\times D_{v}}$, can be written as $\mathbf{w}_{i}=\prod_{j=1}^{i}e^{\mathbf{A}\boldsymbol{\Delta}_{a-1+j}^{\top}}$, representing the cumulative attention weight at step $i$ computed along the scanning path.
Visualization of Activation Maps.
To gain an intuitive and in-depth understanding of SS2D, we further visualize the attention values in $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ corresponding to a specific query patch within foreground objects (referred to as the ‘activation map’). As shown in Figure 5 (b), the activation map of $\mathbf{Q}\mathbf{K}^{\top}$ demonstrates the effectiveness of SS2D in capturing and retaining the traversed information, with all previously scanned tokens in the foreground region being activated. Furthermore, the inclusion of $\mathbf{w}$ leads to activation maps that are more focused on the neighborhood of query patches (Figure 5 (c)), which is consistent with the temporal weighting effect inherent to the formulation of $\mathbf{w}$. Nevertheless, the selective scan mechanism enables VMamba to accumulate the history along the scanning path, thereby facilitating the establishment of long-term dependencies across image patches. This is evident in the sub-figure encircled with a red box (Figure 5 (d)), where patches of the sheep far to the left (scanned in early steps) remain activated. For more visualizations and further discussion, please refer to Appendix D.
Visualization of Effective Receptive Fields.
The Effective Receptive Field (ERF) [41, 11] refers to the region in the input space that contributes to the activation of a certain output unit. We conduct a comparative analysis of the central pixel’s ERF across various visual backbones, both before and after training. The results presented in Figure 6 illustrate that among the models examined, only DeiT, HiViT, and VMamba, demonstrate global ERFs, whereas the others exhibit local ERFs, despite their theoretical potential for global coverage. Moreover, the linear time complexity of VMamba makes it more computationally efficient compared to DeiT and HiViT, which have quadratic costs w.r.t. the number of input patches.
Figure 6: Comparison of Effective Receptive Fields (ERF) [41] between VMamba and other benchmark models. Pixels with higher intensity indicate larger responses regarding the central pixel.
Figure 7: Performance comparison of different scanning patterns.
Diagnostic Study on Selective Scan Patterns.
We compare the proposed scanning pattern (i.e. Cross-Scan) to three benchmark patterns, including unidirectional scanning (Unidi-Scan), bidirectional scanning (Bidi-Scan), and cascade scanning (Cascade-Scan, scanning the data row-wise and column-wise successively). Feature dimensions are adjusted to maintain similar architectural parameters and FLOPs for fair comparison. As illustrated in Figure 7, Cross-Scan outperforms other scanning patterns in both computational efficiency and classification accuracy, highlighting its effectiveness in achieving 2D-Selective-Scan. Removing the DWConv layer, which has been observed to potentially aid the model in learning 2D spatial information, further enhances this advantage. This emphasizes the inherent strength of Cross-Scan in capturing 2D contextual information through its adoption of four-way scanning.
6 Conclusion
This paper presents VMamba, an efficient vision backbone network built with State Space Models (SSMs). VMamba integrates the benefits of selective SSMs from NLP tasks into visual data processing, bridging the gap between ordered 1D scanning and non-sequential 2D traversal through the novel SS2D module. Moreover, we have substantially improved the inference speed of VMamba through a series of architectural and implementation refinements. The effectiveness of the VMamba family has been demonstrated through extensive experiments, and the linear time complexity of VMamba makes it advantageous for downstream tasks with large-resolution inputs.
Limitations.
While VMamba demonstrates promising experimental results, there is still room for this study to be further improved. Previous research has validated the efficacy of unsupervised pre-training on large-scale datasets (e.g., ImageNet-21K). However, the compatibility of existing pre-training methods with SSM-based architectures like VMamba, and the identification of pre-training techniques specifically tailored to such models, remain unexplored. Investigating these aspects could serve as a promising avenue for future research in architectural design. Moreover, limited computational resources have prevented us from exploring VMamba’s architecture at the Large scale, as well as conducting a fine-grained search of hyperparameters to further improve experimental performance.
Appendix A Discretization of State Space Models (SSMs)
In this section, we explore the correlation between the discretized formulation of State Space Models (SSMs) obtained in Sec. 3 and those derived from the zero-order hold (ZOH) method [17], which is frequently used in studies related to SSM.
Recall the discretized formulation of SSMs derived in Sec. 3 as follows,
$\mathbf{h}_{b}=e^{\mathbf{A}\left(\Delta_{a}+...+\Delta_{b-1}\right)}\left(\mathbf{h}_{a}+\sum^{b-1}_{i=a}\mathbf{B}_{i}u_{i}e^{-\mathbf{A}\left(\Delta_{a}+...+\Delta_{i}\right)}\Delta_{i}\right).$ (5)
Let $b=a+1$, then the above equation can be re-written as
$\mathbf{h}_{a+1}$ $=e^{\mathbf{A}\Delta_{a}}\mathbf{h}_{a}+\mathbf{B}_{a}\Delta_{a}u_{a},$ (6)
where $\mathbf{\overline{A_{a}}}\coloneq e^{\mathbf{A}\Delta_{a}}$ is exactly the discretized form of the evolution matrix $\mathbf{A}$ obtained by ZOH, while $\mathbf{\overline{B_{a}}}\coloneq\mathbf{B}_{a}\Delta_{a}$ is basically the first-order Taylor expansion of the discretized $\mathbf{B}$ acquired through ZOH.
Appendix B Derivation of the Recurrence Relation of Selective SSMs
In this section, we derive the recurrence relation of the hidden state in selective SSMs. Given the expression of $h_{b}$ shown in Eq. 5, let us denote $e^{\mathbf{A}\left(\Delta_{a}+...+\Delta_{i-1}\right)}$ as $\mathbf{p_{A,a}^{i}}$, then its recurrence relation can be directly written as
$\mathbf{p_{A,a}^{i}}=e^{\mathbf{A}\Delta_{i-1}}\mathbf{p_{A,a}^{i-1}}.$ (7)
For the second term of Eq. 5, we have
$\mathbf{p_{B,a}^{b}}$ $=e^{\mathbf{A}\left(\Delta_{a}+...+\Delta_{b-1}\right)}\sum^{b-1}_{i=a}\mathbf{B_{i}}u_{i}e^{-\mathbf{A}\left(\Delta_{a}+...+\Delta_{i}\right)}\Delta_{i}$ (8)
$=e^{\mathbf{A}\Delta_{b-1}}\mathbf{p_{B,a}^{b-1}}+\mathbf{B_{b-1}}u_{b-1}\Delta_{b-1}.$ (9)
Therefore, with the associations derived in Eq. 7 and Eq. 8, $\mathbf{h}_{b}=\mathbf{p_{A,a}^{b}}\mathbf{h}_{a}+\mathbf{p_{B,a}^{b}}$ can be efficiently computed in parallel using associative scan algorithms [2, 43, 51], which are supported by numerous modern programming libraries. This approach effectively reduces the overall computational complexity to linear, and VMamba further accelerates the computation by adopting a hardware-aware implementation [17].
Appendix C Details of the relationship between SS2D and Self-attention
In this section, we clarify the relationship between SS2D and the self-attention mechanism commonly employed in existing vision backbone models. Subsequently, visualization results are provided to substantiate our explanation.
Let $T$ denote the length of the sequence with indices from $a$ to $b$, we define the following variables
$\mathbf{V}$ $\coloneqq\left[\mathbf{V_{1}};\ldots;\mathbf{V_{T}}\right]\in\mathbb{R}^{T\times D_{v}},\>\text{where}\>\mathbf{V_{i}}\coloneqq\mathbf{u_{a+i-1}}\odot\boldsymbol{\Delta_{a+i-1}}\in\mathbb{R}^{1\times D_{v}}$ (10)
$\mathbf{K}$ $\coloneqq\left[\mathbf{K_{1}};\ldots;\mathbf{K_{T}}\right]\in\mathbb{R}^{T\times D_{k}},\>\text{where}\>\mathbf{K_{i}}\coloneqq\mathbf{B_{a+i-1}}\in\mathbb{R}^{1\times D_{k}}$ (11)
$\mathbf{Q}$ $\coloneqq\left[\mathbf{Q_{1}};\ldots;\mathbf{Q_{T}}\right]\in\mathbb{R}^{T\times D_{k}},\>\text{where}\>\mathbf{Q_{i}}\coloneqq\mathbf{C_{a+i-1}}\in\mathbb{R}^{1\times D_{k}}$ (12)
$\mathbf{w}$ $\coloneqq\left[\mathbf{w_{1}};\ldots;\mathbf{w_{T}}\right]\in\mathbb{R}^{T\times D_{k}\times D_{v}},\>\text{where}\>\mathbf{w_{i}}\coloneqq\prod_{j=1}^{i}e^{\mathbf{A}\boldsymbol{\Delta_{a-1+j}^{\top}}}\in\mathbb{R}^{D_{k}\times D_{v}}$ (13)
$\mathbf{H}$ $\coloneqq\left[\mathbf{h_{a}};\ldots;\mathbf{h_{b}}\right]\in\mathbb{R}^{T\times D_{k}\times D_{v}},\>\text{where}\>\mathbf{h_{i}}\in\mathbb{R}^{D_{k}\times D_{v}}$ (14)
$\mathbf{Y}$ $\coloneqq\left[\mathbf{y_{a}};\ldots;\mathbf{y_{b}}\right]\in\mathbb{R}^{T\times D_{v}},\>\text{where}\>\mathbf{y_{i}}\in\mathbb{R}^{D_{v}}$ (15)
Note that in practice, the parameter $A$ in Eq. 1 is simplified to $\mathbb{R}^{1\times D_{k}}$. Consequently, $\mathbf{h^{\prime}}(t)=\mathbf{A}\mathbf{h}(t)+\mathbf{B}u(t)$ is simplified to $\mathbf{h^{\prime}}(t)=\mathbf{A}\odot\mathbf{h}(t)+\mathbf{B}u(t)$, which is the reason why $\mathbf{w_{i}}\in\mathbb{R}^{D_{k}\times D_{v}}$.
Based on these notations, the discretized solution of time-varying SSMs (Eq. 5) can be written as
$\mathbf{h_{b}}=\mathbf{w_{T}}\odot\mathbf{h_{a}}+\sum_{i=1}^{T}\frac{\mathbf{w_{T}}}{\mathbf{w_{i}}}\odot\left(\mathbf{K_{i}}^{\top}\mathbf{V_{i}}\right),$ (16)
where $\odot$ denotes the element-wise product between matrices, and the division is also elements-wise.
Based on the expression of the hidden state $\mathbf{h_{b}}$, the first term of the output of SSM, i.e., $\mathbf{y_{b}}$, can be computed by
$\mathbf{y_{b}}$ $=\mathbf{Q_{T}}\mathbf{h_{b}}$ (17)
$=\mathbf{Q_{T}}\left(\mathbf{w_{T}}\odot\mathbf{h_{a}}\right)+\mathbf{Q_{T}}\sum_{i=1}^{T}\frac{\mathbf{w_{T}}}{\mathbf{w_{i}}}\odot\left(\mathbf{K_{i}}^{\top}\mathbf{V_{i}}\right).$ (18)
[…]
$\mathbf{y_{b}}^{(j)}=\left(\mathbf{Q_{T}}\odot\mathbf{w_{T}}^{(j)}\right){\mathbf{h_{a}}^{(j)}}+\sum_{i=1}^{T}\left(\frac{\mathbf{Q_{T}}\odot\mathbf{w_{T}}^{(j)}}{\mathbf{w_{i}}^{(j)}}\mathbf{K_{i}}^{\top}\right)\odot\mathbf{V_{i}}^{(j)}.$ (19)
[…]
$\mathbf{Y}^{(j)}=\left(\mathbf{Q}\odot\mathbf{w}^{(j)}\right){\mathbf{h_{a}}^{(j)}}+\left[\left(\mathbf{Q}\odot\mathbf{w}^{(j)}\right)\left(\frac{\mathbf{K}}{\mathbf{w}^{(j)}}\right)^{\top}\odot\mathbf{M}\right]\mathbf{V}^{(j)},$ (20)
where $\mathbf{M}\coloneqq\texttt{tril}(T,T)\in\{0,1\}^{T\times T}$ denotes the temporal mask matrix with the lower triangular portion of a $T\times T$ matrix set to 1 and elsewhere 0. It is evident that how matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are multiplied in Eq. 20 closely resembles the process in the self-attention module of Vision Transformers. Moreover, if $\mathbf{w}$ is in shape $(T,D_{k})$ rather than $(T,D_{k},D_{v})$, then Eq. 17 and Eq. 20 reduces to the form of Gated Linear Attention (GLA) [67], which indicates that GLA is also a special case of Mamba.
Appendix D Visualization of Attention and Activation Maps
In the preceding subsection, we illustrated how the computational process of selective SSMs shares similarities with self-attention mechanisms, allowing us to delve into the internal mechanism of SS2D through the visualization of its weight matrices.
Given the input image as in Figure 8 (a), illustrations of four scanning paths in SS2D are shown in Figure 8 (d), the visualization of corresponding attention maps by calculating $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ are shown in Figure 8 (e) and Figure 8 (g) respectively. These results underscore the effectiveness of the proposed scanning approach (i.e., Cross-Scan) in capturing and retaining the traversed information, as each row in a single attention map corresponds to the attention between the current patch and all previously scanned foreground tokens impartially. Additionally, in Figure 8 (f), we showcase the transformed activation maps, where the pixel order corresponds to that of the first route, traversing the image row-wise from the upper-left to the bottom-right.
By rearranging the diagonal elements of the obtained attention map in the image space, we derive the visualization results illustrated in Figure 8 (b) and Figure 8 (c) corresponding to $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ respectively. These maps illustrate the effectiveness of VMamba in accurately distinguishing between foreground and background pixels within an image.
Moreover, given a certain selected patch as the query, we visualize the corresponding ‘activation map’ by reshaping the associated row in the attention map (computed by $\mathbf{Q}\mathbf{K}^{\top}$ or $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$) This reflects the attention score between the query patch and all previously scanned patches. To get the complete visualization for a query patch, we collect and combine the activation maps from all four scanning paths in SS2D. Visualization results of the activation map for both $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ are shown in Figure 9. Moreover, we also visualize the diagonal elements of attention maps computed by $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$, where all foreground objects are effectively highlighted and separated from the background.
Figure 8: Illustration of the attention maps obtained by SS2D.
Figure 9: Illustration of activation maps for the query patch (marked with a red star).
Appendix E Detailed Experiment Settings
Network Architecture.
The architectural specifications of Vanilla-VMamba are outlined in Table 3, and detailed configurations of the VMamba series are provided in Table 4. The Vanilla-VMamba series are built with the vanilla VSS Block, which includes a multiplicative branch with no feed-forward network (FFN) layer. In contrast, the VSS Block in the VMamba series eliminates the multiplicative branch and introduces FFN layers. Additionally, we provide alternative architectures for VMamba at Small and Base scales, referred to as VMamba-S[$s1l20$] and VMamba-B[$s1l20$], respectively. The notation ‘$sxly$’ indicates that the ssm-ratio is set to $x$ and the number of layers in stage 3 is set to $y$. Consequently, the versions presented in Table 1 can also be referred to as VMamba-S[$s2l15$] and VMamba-B[$s2l15$].
Experiment Setting.
The hyper-parameters for training VMamba on ImageNet are inherited from Swin [37], except for the parameters related to drop_path_rate and the exponential moving average (EMA) technique. Specifically, VMamba-T/S/B models are trained from scratch for 300 epochs, with a 20-epoch warm-up period, using a batch size of 1024. The training process utilizes the AdamW optimizer [39] with betas set to $(0.9,0.999)$, a momentum of $0.9$, an initial learning rate of $1\times 10^{-3}$, a weight decay of $0.05$, and a cosine decay learning rate scheduler. Additional techniques such as label smoothing (0.1) and EMA (decay ratio of 0.9999) are also applied. The drop_path_ratio is set to 0.2 for Vanilla-VMamba-T, VMamba-T, 0.3 for Vanilla-VMamba-S, VMamba-S[$s2l15$], VMamba-S[$s1l20$], 0.6 for Vanilla-VMamba-B, VMamba-B[$s2l15$] and 0.5 for VMamba-B[$s1l20$]. No other training techniques are employed.
Throughput Evaluation.
Detailed performance comparisons with various models are presented in Table 6. Throughput (referred to as TP.) was assessed on an A100 GPU paired with an AMD EPYC 7542 CPU, utilizing the toolkit provided by [64]. Following the protocol outlined in [37], we set the batch size to 128. The training throughput (referred to as Train TP.) is tested on the same device with mix-resolution, excluding the time consumption of optimizers. The batch size for measuring the training throughput is set to 128 as well.
Accelerating VMamba.
Table 5 provides detailed configurations of intermediate variants in the acceleration process from Vanilla-VMamba-T to VMamba-T.
Evolution of ERF.
We further generate the effective receptive field (ERF) maps throughout the training process for Vanilla-VMamba-T. These maps intuitively illustrate how VMamba’s pattern of ERF evolves from being predominantly local to predominantly global, epoch by epoch.
Table 3: Architectural overview of the Vanilla-VMamba series. Down-sampling is executed through patch merging [37] operations in stages 1, 2, and 3. The term Linear refers to a linear layer. The DWConv denotes a depth-wise convolution [24] operation. The proposed 2D-selective-scan is labeled as SS2D.
layer name output size Vanilla-VMamba-T Vanilla-VMamba-S Vanilla-VMamba-B
stem 112$\times$112 conv 4$\times$4, 96, stride 4 conv 4$\times$4, 96, stride 4 conv 4$\times$4, 128, stride 4
stage 1 56$\times$56
vanilla VSSBLock vanilla VSSBLock vanilla VSSBLock
$\left[\begin{array}[]{c}\text{{Linear} 96 $\rightarrow 2\times$96}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$96}\\[-1.00006pt] \text{{SS2D}, dim $2\times$96}\\[-1.00006pt] \text{{Linear} 2$\times$96 $\rightarrow$ 96}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$96 $\rightarrow$ 96}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 96 $\rightarrow 2\times$96}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$96}\\[-1.00006pt] \text{{SS2D}, dim $2\times$96}\\[-1.00006pt] \text{{Linear} 2$\times$96 $\rightarrow$ 96}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$96 $\rightarrow$ 96}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 128 $\rightarrow 2\times$128}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$128}\\[-1.00006pt] \text{{SS2D}, dim $2\times$128}\\[-1.00006pt] \text{{Linear} 2$\times$128 $\rightarrow$ 128}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$128 $\rightarrow$ 128}\\[-1.00006pt] \end{array}\right]$$\times$2
patch merging $\rightarrow 192$ patch merging $\rightarrow 192$ patch merging $\rightarrow 256$
stage 2 28$\times$28
vanilla VSSBLock vanilla VSSBLock vanilla VSSBLock
$\left[\begin{array}[]{c}\text{{Linear} 192 $\rightarrow 2\times$192}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$192}\\[-1.00006pt] \text{{SS2D}, dim $2\times$192}\\[-1.00006pt] \text{{Linear} 2$\times$192 $\rightarrow$ 192}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$192 $\rightarrow$ 192}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 192 $\rightarrow 2\times$192}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$192}\\[-1.00006pt] \text{{SS2D}, dim $2\times$192}\\[-1.00006pt] \text{{Linear} 2$\times$192 $\rightarrow$ 192}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$192 $\rightarrow$ 192}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 256 $\rightarrow 2\times$256}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$256}\\[-1.00006pt] \text{{SS2D}, dim $2\times$256}\\[-1.00006pt] \text{{Linear} 2$\times$256 $\rightarrow$ 256}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$256 $\rightarrow$ 256}\\[-1.00006pt] \end{array}\right]$$\times$2
patch merging $\rightarrow 384$ patch merging $\rightarrow 384$ patch merging $\rightarrow 512$
stage 3 14$\times$14
vanilla VSSBLock vanilla VSSBLock vanilla VSSBLock
$\left[\begin{array}[]{c}\text{{Linear} 384 $\rightarrow 2\times$384}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$384}\\[-1.00006pt] \text{{SS2D}, dim $2\times$384}\\[-1.00006pt] \text{{Linear} 2$\times$384 $\rightarrow$ 384}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$384 $\rightarrow$ 384}\\[-1.00006pt] \end{array}\right]$$\times$9 $\left[\begin{array}[]{c}\text{{Linear} 384 $\rightarrow 2\times$384}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$384}\\[-1.00006pt] \text{{SS2D}, dim $2\times$384}\\[-1.00006pt] \text{{Linear} 2$\times$384 $\rightarrow$ 384}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$384 $\rightarrow$ 384}\\[-1.00006pt] \end{array}\right]$$\times$27 $\left[\begin{array}[]{c}\text{{Linear} 512 $\rightarrow 2\times$512}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$512}\\[-1.00006pt] \text{{SS2D}, dim $2\times$512}\\[-1.00006pt] \text{{Linear} 2$\times$512 $\rightarrow$ 512}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$512 $\rightarrow$ 512}\\[-1.00006pt] \end{array}\right]$$\times$27
patch merging $\rightarrow 768$ patch merging $\rightarrow 768$ patch merging $\rightarrow 1024$
stage 4 7$\times$7
vanilla VSSBLock vanilla VSSBLock vanilla VSSBLock
$\left[\begin{array}[]{c}\text{{Linear} 768 $\rightarrow 2\times$768}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$768}\\[-1.00006pt] \text{{SS2D}, dim $2\times$768}\\[-1.00006pt] \text{{Linear} 2$\times$768 $\rightarrow$ 768}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$768 $\rightarrow$ 768}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 768 $\rightarrow 2\times$768}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$768}\\[-1.00006pt] \text{{SS2D}, dim $2\times$768}\\[-1.00006pt] \text{{Linear} 2$\times$768 $\rightarrow$ 768}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$768 $\rightarrow$ 768}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 1024 $\rightarrow 2\times$1024}\\[-1.00006pt] \text{{DWConv} 3$\times$3, $2\times$1024}\\[-1.00006pt] \text{{SS2D}, dim $2\times$1024}\\[-1.00006pt] \text{{Linear} 2$\times$1024 $\rightarrow$ 1024}\\[-1.00006pt] \text{{Multiplicative}}\\[-1.00006pt] \text{{Linear} 2$\times$1024 $\rightarrow$ 1024}\\[-1.00006pt] \end{array}\right]$$\times$2
1$\times$1 average pool, 1000-d fc, softmax
Param. (M) 22.9 44.4 76.3
FLOPs 5.63×$\times 10^{9}$ 11.23×$\times 10^{9}$ 18.02×$\times 10^{9}$
Table 4: Architectural overview of the VMamba series.
layer name output size VMamba-T VMamba-S VMamba-B
stem 112$\times$112 conv 3$\times$3 stride 2, LayerNorm, GeLU, conv 3$\times$3 stride 2, LayerNorm
stage 1 56$\times$56
VSSBLock(ssm-ratio=1, mlp-ratio=4) VSSBlock(ssm-ratio=2, mlp-ratio=4) VSSBLock(ssm-ratio=2, mlp-ratio=4)
$\left[\begin{array}[]{c}\text{{Linear} 96 $\rightarrow$ {ssm-ratio} $\times$96}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$96}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$96}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$96 $\rightarrow$ 96}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$96}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 96 $\rightarrow$ {ssm-ratio} $\times$96}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$96}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$96}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$96 $\rightarrow$ 96}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$96}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 128 $\rightarrow$ {ssm-ratio} $\times$128}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$128}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$128}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$128 $\rightarrow$ 128}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$128}\\[-1.00006pt] \end{array}\right]$$\times$2
conv 3$\times$3 stride 2, LayerNorm
stage 2 28$\times$28
VSSBLock(ssm-ratio=1, mlp-ratio=4) VSSBlock(ssm-ratio=2, mlp-ratio=4) VSSBLock(ssm-ratio=2, mlp-ratio=4)
$\left[\begin{array}[]{c}\text{{Linear} 192 $\rightarrow$ {ssm-ratio} $\times$192}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$192}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$192}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$192 $\rightarrow$ 192}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$192}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 192 $\rightarrow$ {ssm-ratio} $\times$192}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$192}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$192}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$192 $\rightarrow$ 192}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$192}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 256 $\rightarrow$ {ssm-ratio} $\times$256}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$256}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$256}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$256 $\rightarrow$ 256}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$256}\\[-1.00006pt] \end{array}\right]$$\times$2
conv 3$\times$3 stride 2, LayerNorm
stage 3 14$\times$14
VSSBLock(ssm-ratio=1, mlp-ratio=4) VSSBlock(ssm-ratio=2, mlp-ratio=4) VSSBLock(ssm-ratio=2, mlp-ratio=4)
$\left[\begin{array}[]{c}\text{{Linear} 384 $\rightarrow$ {ssm-ratio} $\times$384}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$384}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$384}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$384 $\rightarrow$ 384}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$384}\\[-1.00006pt] \end{array}\right]$$\times$8 $\left[\begin{array}[]{c}\text{{Linear} 384 $\rightarrow$ {ssm-ratio} $\times$384}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$384}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$384}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$384 $\rightarrow$ 384}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$384}\\[-1.00006pt] \end{array}\right]$$\times$15 $\left[\begin{array}[]{c}\text{{Linear} 512 $\rightarrow$ {ssm-ratio} $\times$512}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$512}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$512}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$512 $\rightarrow$ 512}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$512}\\[-1.00006pt] \end{array}\right]$$\times$15
conv 3$\times$3 stride 2, LayerNorm
stage 4 7$\times$7
VSSBLock(ssm-ratio=1, mlp-ratio=4) VSSBlock(ssm-ratio=2, mlp-ratio=4) VSSBLock(ssm-ratio=2, mlp-ratio=4)
$\left[\begin{array}[]{c}\text{{Linear} 768 $\rightarrow$ {ssm-ratio} $\times$768}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$768}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$768}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$768 $\rightarrow$ 768}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$768}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 768 $\rightarrow$ {ssm-ratio} $\times$768}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$768}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$768}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$768 $\rightarrow$ 768}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$768}\\[-1.00006pt] \end{array}\right]$$\times$2 $\left[\begin{array}[]{c}\text{{Linear} 1024 $\rightarrow$ {ssm-ratio} $\times$1024}\\[-1.00006pt] \text{{DWConv} 3$\times$3, {ssm-ratio} $\times$1024}\\[-1.00006pt] \text{{SS2D}, dim {ssm-ratio} $\times$1024}\\[-1.00006pt] \text{{Linear} {ssm-ratio} $\times$1024 $\rightarrow$ 1024}\\[-1.00006pt] \text{{FFN} {mlp-ratio} $\times$1024}\\[-1.00006pt] \end{array}\right]$$\times$2
1$\times$1 average pool, 1000-d fc, softmax
Param. (M) 30.2 50.1 88.6
FLOPs 4.91×$\times 10^{9}$ 8.72×$\times 10^{9}$ 15.36×$\times 10^{9}$
Table 5: Details of accelerating VMamba.
  Model d_state ssm-ratio DWConv multiculative layers FFN Params FLOPs TP. Train TP. Top-1
branch numbers (M) (G) (img/s) (img/s) (%)
Vanilla-VMamba-T 16 2.0 [2,2,9,2] 22.9M 5.63G 426 138 82.17
Step(a) 16 2.0 [2,2,9,2] 22.9M 5.63G 467 165 82.17
Step(b) 16 2.0 [2,2,9,2] 22.9M 5.63G 464 184 82.17
Step(c) 16 2.0 [2,2,9,2] 22.9M 5.63G 638 195 82.17
Step(d) 16 2.0 [2,2,2,2] 29.0M 5.63G 813 248 81.65
Step(e) 16 1.0 [2,2,5,2] 26.2M 4.86G 1179 360 82.17
Step(f) 1 2.0 [2,2,5,2] 30.7M 4.86G 1340 464 82.49
Step(g) 1 1.0 [2,2,8,2] 30.2M 4.91G 1686 571 82.60
Table 6: Performance comparison on ImageNet-1K with image size 224. indicates that Vim is trained only in float32 in practical, in which case the train throughput is 232. [71].
  Model Image Params FLOPs TP. Train TP. Top-1
Size (M) (G) (img/s) (img/s) (%)
DeiT-S [58] 2242 22M 4.6G 1761 2404 79.8
DeiT-B [58] 2242 86M 17.5G 503 1032 81.8
ConvNeXt-T [38] 2242 29M 4.5G 1198 702 82.1
ConvNeXt-S [38] 2242 50M 8.7G 684 445 83.1
ConvNeXt-B [38] 2242 89M 15.4G 436 334 83.8
HiViT-T [68] 2242 19M 4.6G 1393 1304 82.1
HiViT-S [68] 2242 38M 9.1G 712 698 83.5
HiViT-B [68] 2242 66M 15.8G 456 544 83.8
Swin-T [37] 2242 28M 4.5G 1244 987 81.3
Swin-S [37] 2242 50M 8.7G 718 642 83.0
Swin-B [37] 2242 88M 15.5G 458 496 83.5
XCiT-S12/16 2242 26M 4.9G 1283 935 82.0
XCiT-S24/16 2242 48M 9.2G 671 509 82.6
XCiT-M24/16 2242 84M 16.2G 423 385 82.7
S4ND-ConvNeXt-T [45] 2242 30M - 683 369 82.2
S4ND-ViT-B [45] 2242 89M - 398 400 80.4
Vim-S [71] 2242 26M - 811 344 80.5
Vanilla-VMamba-T 2242 23M 5.6G 638 195 82.2
Vanilla-VMamba-S 2242 44M 11.2G 359 111 83.5
Vanilla-VMamba-B 2242 76M 18.0G 268 84 83.7
VMamba-T 2242 30M 4.9G 1686 571 82.6
VMamba-S[$s2l15$] 2242 50M 8.7G 877 314 83.6
VMamba-B[$s2l15$] 2242 89M 15.4G 646 247 83.9
VMamba-S[$s1l20$] 2242 49M 8.6G 1106 390 83.3
VMamba-B[$s1l20$] 2242 87M 15.2G 827 313 83.8
Figure 10: Illustration of ERF maps throughout the training process of Vanilla-VMamba-T (with EMA).
Appendix F Performance of the VMamba Family on Downstream Tasks
In this section, we present the experimental results of Vanilla-VMamba and VMamba on the MSCOCO and ADE20k datasets. The results are summarized in Table 7 and Table 8, respectively.
For object detection and instance segmentation, we adhere to the protocol outlined by Swin [37] and construct our models using the mmdetection framework [3]. Specifically, we utilize the AdamW optimizer [39] and fine-tune the classification models pre-trained on ImageNet-1K for both 12 and 36 epochs. The learning rate is initialized with $1\times 10^{-4}$ and is decreased by a factor of 10 at the 9-th and 11-th epoch. We incorporate multi-scale training and random flipping with a batch size of 16, in line with established practices for object detection evaluations.
For semantic segmentation, we follow Swin [37] and construct a UperHead [65] network on top of the pre-trained model using the MMSegmentation library [4]. We employ the AdamW optimizer [39] and set the learning rate to $6\times 10^{-5}$. The fine-tuning process spans a total of $160k$ iterations with a batch size of 16. The default input resolution is $512\times 512$.
Table 7: Object detection and instance segmentation results on COCO dataset. The FLOPs are calculated using inputs of size $1280\times 800$. Here, $AP^{b}$ and $AP^{m}$ denote box AP and mask AP, respectively. "1×1$\times$1 ×" indicates models fine-tuned for 12 epochs, while "3×3$\times$3 ×MS" signifies the utilization of multi-scale training for $36$ epochs.
Mask R-CNN 1$\times$ schedule
[…]
Swin-T 42.7 65.2 46.8 39.3 62.2 42.2 48M 267G
ConvNeXt-T 44.2 66.6 48.3 40.1 63.3 42.8 48M 262G
Vanilla-VMamba-T 46.5 68.5 50.7 42.1 65.5 45.3 42M 286G
VMamba-T 47.3 69.3 52.0 42.7 66.4 45.9 50M 271G
Swin-S 44.8 66.6 48.9 40.9 63.4 44.2 69M 354G
ConvNeXt-S 45.4 67.9 50.0 41.8 65.2 45.1 70M 348G
Vanilla-VMamba-S 48.2 69.7 52.5 43.0 66.6 46.4 64M 400G
VMamba-S 48.7 70.0 53.4 43.7 67.3 47.0 70M 349G
Swin-B 46.9 42.3 107M 496G
ConvNeXt-B 47.0 69.4 51.7 42.7 66.3 46.0 108M 486G
Vanilla-VMamba-B 48.6 70.0 53.1 43.3 67.1 46.7 96M 540G
VMamba-B 49.2 71.4 54.0 44.1 68.3 47.7 108M 485G
Mask R-CNN 3$\times$ MS schedule
Swin-T 46.0 68.1 50.3 41.6 65.1 44.9 48M 267G
ConvNeXt-T 46.2 67.9 50.8 41.7 65.0 44.9 48M 262G
Vanilla-VMamba-T 48.5 70.0 52.7 43.2 66.9 46.4 42M 286G
VMamba-T 48.8 70.4 53.5 43.7 67.4 47.0 50M 271G
Swin-S 48.2 69.8 52.8 43.2 67.0 46.1 69M 354G
ConvNeXt-S 47.9 70.0 52.7 42.9 66.9 46.2 70M 348G
Vanilla-VMamba-S 49.7 70.4 54.2 44.0 67.6 47.3 64M 400G
VMamba-S 49.9 70.9 54.7 44.2 68.2 47.7 70M 349G
Table 8: Semantic segmentation results on ADE20K using UperNet [65]. We evaluate the performance of semantic segmentation on the ADE20K dataset with UperNet [65]. The FLOPs are calculated with input sizes of $512\times 2048$. "SS" and "MS" denote single-scale and multi-scale testing, respectively.
method crop size mIoU (SS) mIoU (MS) Params FLOPs
Swin-T $512^{2}$ 44.5 45.8 60M 945G
ConvNeXt-T $512^{2}$ 46.0 46.7 60M 939G
Vanilla-VMamba-T $512^{2}$ 47.3 48.3 55M 964G
VMamba-T $512^{2}$ 48.0 48.8 62M 949G
Swin-S $512^{2}$ 47.6 49.5 81M 1039G
ConvNeXt-S $512^{2}$ 48.7 49.6 82M 1027G
Vanilla-VMamba-S $512^{2}$ 49.5 50.5 76M 1081G
VMamba-S $512^{2}$ 50.6 51.2 82M 1028G
Swin-B $512^{2}$ 48.1 49.7 121M 1188G
ConvNeXt-B $512^{2}$ 49.1 49.9 122M 1170G
Vanilla-VMamba-B $512^{2}$ 50.0 51.3 110M 1226G
VMamba-B $512^{2}$ 51.0 51.6 122M 1170G
Appendix G Details of VMamba’s Scale-Up Experiments
Given Mamba’s exceptional ability in efficient long sequence modeling, we conduct experiments to assess whether VMamba inherits this characteristic. We evaluate the computational efficiency and classification accuracy of VMamba with progressively larger input spatial resolutions. Specifically, following the protocol in XCiT [1], we apply VMamba trained on $224\times 224$ input to images with resolutions ranging from $288\times 288$ to $768\times 768$. We measure the generalization performance in terms of the number of parameters, FLOPs, throughput during both training and inference, and the top-1 classification accuracy on ImageNet-1K. We also conduct experiments under the ‘linear tuning’ setting, where only the header network, consisting of a single linear module, is fine-tuned from random initialization using features extracted by those backbone models.
According to the results summarized in Table 9, VMamba demonstrates the most stable performance across (i.e., modest performance drop) different input image sizes, achieving a top-1 classification accuracy of $74.7\%$ without fine-tuning ($79.2\%$ with linear tuning), while maintaining a relatively high throughput of 149 images per second at an input resolution of $768\times 768$. In comparison, Swin [37] achieves the second-highest performance with a top-1 accuracy of $73.1\%$ without fine-tuning ($77.5\%$ under linear tuning) at the same input size, when using scaled window sizes (set as the resolution divided by $32$). However, its throughput significantly drops to $53$ images per second. Furthermore, ConvNeXt [38] maintains a relatively high inference speed (i.e., a throughput of $103$ images per second) with the largest input resolution. However, its classification accuracy drops to $69.5\%$ when directly tested on images of size $768\times 768$, indicating its limited adaptability w.r.t. input images with large spatial resolutions. The performance of Deit-S also drops dramatically, mainly due to the use of interpolation in the absolute positional embedding.
The performance variations across different resolutions for various models are more intuitively illustrated in Figure 11. Notably, VMamba displays a linear increase in computational complexity (Figure 4 (b)), as measured by FLOPs, which is comparable to CNN-based architectures. This finding aligns with the theoretical conclusions drawn from selective SSMs [17].
[…]
Model Image Param. FLOPs TP. Train TP. Top-1 Top-1 acc. (%)
Size (M) (G) (img/s) (img/s) acc. (%) (w/ linear tuning)
SSM-Based
VMamba-Tiny 2242 30M 4.91G 1490 418 82.60 82.64
2882 30M 8.11G 947 303 82.95 83.03
3842 30M 14.41G 566 187 82.41 82.77
5122 30M 25.63G 340 121 80.92 81.88
6402 30M 40.04G 214 75 78.60 80.62
7682 30M 57.66G 149 53 74.66 79.22
VMamba-Tiny[$s2l5$] 2242 31M 4.86G 1227 399 82.49 82.52
2882 31M 8.03G 761 255 82.81 82.93
3842 31M 14.27G 452 155 82.51 82.74
5122 31M 25.38G 272 100 81.07 82.02
6402 31M 39.65G 170 60 79.30 81.02
7682 31M 57.09G 117 42 76.06 79.69
Vanilla-VMamba-Tiny 2242 23M 5.63G 628 189 82.17 82.09
2882 23M 9.30G 390 117 82.74 82.76
3842 23M 16.53G 212 65 82.40 82.72
5122 23M 29.39G 138 53 81.05 81.97
6402 23M 45.93G 87 27 78.79 80.71
7682 23M 66.14G 52 18 75.09 79.12
Transformer-Based
Swin-Tiny 2242 28M 4.51G 1142 769 81.19 81.18
2882 28M 7.60G 638 489 81.46 81.62
3842 28M 14.05G 316 268 80.67 81.12
5122 28M 26.65G 176 131 78.97 80.21
6402 28M 45.00G 88 68 76.55 78.89
7682 29M 70.72G 53 38 73.06 77.54
XCiT-S12/16 2242 26M 4.87G 1127 505 81.87 81.89
2882 26M 8.05G 724 462 82.44 82.44
3842 26M 14.31G 425 308 81.84 82.21
5122 26M 25.44G 244 185 79.80 80.92
6402 26M 39.75G 158 122 76.84 79.00
7682 26M 57.24G 111 87 72.52 76.92
HiViT-Tiny 2242 19M 4.60G 1261 1041 81.92 81.85
2882 19M 7.93G 750 614 82.45 82.42
3842 19M 15.21G 388 333 81.51 81.91
5122 20M 30.56G 186 150 79.30 80.49
6402 20M 54.83G 93 71 76.09 78.58
7682 20M 91.41G 55 37 71.38 76.47
DeiT-Small 2242 22M 4.61G 1573 1306 80.69 80.40
2882 22M 7.99G 914 1124 80.80 80.63
3842 22M 15.52G 502 697 78.87 79.54
5122 22M 31.80G 261 387 74.21 76.91
6402 23M 58.17G 149 244 68.04 73.31
7682 23M 98.70G 90 156 60.98 69.62
ConvNet-Based
ConvNeXt-Tiny 2242 29M 4.47G 1107 614 82.05 81.95
2882 29M 7.38G 696 403 82.23 82.30
3842 29M 13.12G 402 240 81.05 81.78
5122 29M 23.33G 226 140 78.03 80.37
6402 29M 36.45G 147 90 74.27 78.77
7682 29M 52.49G 103 63 69.50 76.89
Figure 11: Throughtput and Top-1 accuracy change on ImageNet-1K w.r.t resolution rises. Swin-T denotes Swin-T tested with scaled window sizes.
Appendix H Ablation Study
H.1 Influence of the Scanning Pattern
In the main submission, we validate the effectiveness of the proposed scanning pattern (referred to as Cross-Scan) in SS2D by comparing to three alternative image traversal approaches, i.e., Unidi-Scan, Bidi-Scan, and Cascade-Scan (Figure 12). Notably, as Unidi-Scan, Bidi-Scan, and Cross-Scan are all implemented in Triton, they exhibit little difference in throughput. Nevertheless, the findings in Table 10 indicate that Cross-Scan exhibits superior data modeling capacity, evident from its higher classification accuracy. This advantage is likely attributed to the two-dimensional prior introduced by the four-way scanning design. Nevertheless, the practical implementation of Cascade-Scan is notably constrained by its relatively slow computational pace. This limitation primarily stems from the inadequate compatibility between selective scanning and high-dimensional data, exacerbated by the multi-step scanning procedure.
Figure 13 indirectly demonstrates that among the analyzed scanning methods, only Bidi-Scan, Cascade-Scan, and Cross-Scan showcase global ERFs. Moreover, only Cross-Scan and Cascade-Scan exhibit two-dimensional (2D) priors. It is also noteworthy that DWConv [24] plays a significant role in establishing 2D priors, thereby contributing to the formation of global ERFs.
Figure 12: Illustration of different scanning patterns for selective scan.
Table 10: The performance of VMamba-T with different scanning patterns.
  Model Params FLOPs TP. Train TP. Top-1
(M) (G) (img/s) (img/s) (%)
VMamba w/ dwconv
Unidi-Scan 30.2M 4.91G 1682 571 82.26
Bidi-Scan 30.2M 4.91G 1687 572 82.49
Cascade-Scan 30.2M 4.91G 998 308 82.42
Cross-Scan 30.2M 4.91G 1686 571 82.60
VMamba w/o dwconv
Unidi-Scan 30.2M 4.89G 1716 578 80.88
Bidi-Scan 30.2M 4.89G 1719 578 81.80
Cascade-Scan 30.2M 4.90G 1007 309 81.71
Cross-Scan 30.2M 4.89G 1717 577 82.25
Figure 13: The visualization of ERF for models with different scanning patterns.
H.2 Influence of the Initialization Approach
In our study, we adopted the initialization scheme initially proposed for the SS2D block in S4D [19]. Therefore, it is necessary to investigate the contribution of this initialization method to the effectiveness of VMamba. To delve deeper into this matter, we replaced the default initialization with two alternative methods: random initialization and zero initialization.
For both initialization methods, we set the parameter $\mathbf{D}$ in equation 1 to a vector of all ones, mimicking a basic skip connection (thus we have $\mathbf{y}=\mathbf{C}\mathbf{h}+\mathbf{D}\mathbf{u}$). Additionally, the weights and biases associated with the transformation from low rank to the dimension $D_{v}$ (which matches the input size), are initialized as random vectors, in contrast to Mamba [17], which employs a more sophisticated approach for initialization.
The main distinction between random and zero initialization lies in the parameter $A$ in equation 5, which is typically initialized as a HiPPO matrix in both Mamba [17, 19] and our implementation of VMamba. Given that we selected the hyper-parameter d_state to be 1, the Mamba initialization for $log(A)$ can be simplified to all zeros, which aligns with zero initialization. In contrast, random initialization assigns a random vector to $log(A)$. We choose to initialize $log(A)$ rather than $A$ directly to ensure that $A$ remains close to the all-ones matrix when the network parameters are near zero, which empirically helps enhance the stability of the training process.
The experiment results in Table 11 indicate that, at least for image classification with SS2D blocks, the model’s performance is not significantly affected by the initialization method. Therefore, within this context, the sophisticated initialization method employed in Mamba [17] can be substituted with a simpler, more straightforward approach. We also visualize the ERF maps of models trained with different initialization methods (see Figure 14), which intuitively reflect the robustness of SS2D across various initialization schemes.
Table 11: The performance of VMamba-T with different initialization.
  initialization Params FLOPs TP. Train TP. Top-1
(M) (G) (img/s) (img/s) acc. (%)
mamba 30.2 4.91 1686 571 82.60
rand 30.2 4.91 1682 570 82.58
zero 30.2 4.91 1683 570 82.67
Figure 14: The visualization of ERF of VMamba with different initialization.
H.3 Influence of the d_state Parameter
Throughout this study, we primarily set the value of d_state to 1 to optimize the computational speed of VMamba. To further explore the impact of d_state on the model’s performance, we conduct a series of experiments.
As shown in Table 12, with all other hyper-parameters frozen, we increase d_state from 1 to 4. This leads to a minor improvement in performance but a substantial decrease in throughput, indicating a significant negative influence on the computational efficiency of VMamba. However, increasing d_state to 8, while reducing ssm-ratio to maintain computational complexity, leads to improved accuracy. However, when d_state is further increased to 16, with ssm-ratio set to 1, performance drops.
These findings suggest that modest increases in d_state may not necessarily lead to improved performance. Instead, selecting the optimal combination of d_state and ssm-ratio is crucial to achieving a good trade-off between inference speed and performance.
Table 12: The performance of VMamba-T with different d_state.
  d_state ssm-ratio Params FLOPs TP. Train TP. Top-1
(M) (G) (img/s) (img/s) acc. (%)
1 2.0 30.7 4.86 1340 464 82.49
2 2.0 30.8 4.98 1269 432 82.50
4 2.0 31.0 5.22 1147 382 82.51
8 1.5 28.6 5.04 1148 365 82.69
16 1.0 26.3 4.87 1164 358 82.31
H.4 Influence of ssm-ratio, mlp-ratio, and layer numbers
In this section, we investigate the trade-off between ssm-ratio, layer numbers, and mlp-ratio.
Experiment results shown in Table 13 suggest that reducing ssm-ratio leads to a notable decrease in performance but significantly enhances the inference speed. Conversely, increasing layer numbers enhances the performance while slowing down the model.
As the hyper-parameter ssm-ratio stands for the dimension employed by the SS2D module, the trade-off between ssm-ratio and layer numbers can be interpreted as a balance between channel-mixing and token-mixing [57]. Furthermore, we reduce mlp-ratio from 4.0 to 2.0 and progressively increase ssm-ratio to maintain constant FLOPs, as shown in Table 14. The results presented in Tables 13 and 14 underscore the importance of an optimal combination of ssm-ratio, mlp-ratio, and layer numbers for constructing a model that achieves a balance between effectiveness and efficiency.
Table 13: The performance of VMamba-T under different combination of ssm-ratio and layer numbers.
  ssm-ratio layer Params FLOPs TP. Train TP. Top-1
numbers (M) (G) (img/s) (img/s) acc. (%)
2.0 [2,2,5,2] 30.7 4.86 1340 464 82.49
1.0 [2,2,5,2] 25.6 3.98 1942 647 81.87
1.0 [2,2,8,2] 30.2 4.91 1686 571 82.60
Table 14: The performance of VMamba under different combination of ssm-ratio and mlp-ratio.
  mlp-ratio ssm-ratio Params FLOPs TP. Train TP. Top-1
(M) (G) (img/s) (img/s) acc. (%)
4.0 1.0 30.2 4.91 1686 571 82.60
3.0 1.5 28.5 4.65 1419 473 82.75
2.0 2.5 29.9 4.95 1075 361 82.86
H.5 Influence of the Activation Function
In VMamba, the SiLU [14] activation function is utilized to build the SS2D block. However, experiment results in Table 15 reveal that VMamba maintains robustness across different activation functions. This implies that the selection of the activation function does not substantially affect the model’s performance. Therefore, there is flexibility in choosing an appropriate activation function based on computational constraints or other preferences.
Table 15: The performance of VMamba-T with different activation functions in SS2D.
  activation Params FLOPs TP. Train TP. Top-1
(M) (G) (img/s) (img/s) acc. (%)
SiLU 30.2 4.91 1686 571 82.60
GELU 30.2 4.91 1680 570 82.53
ReLU 30.2 4.91 1684 577 82.65